# Federated Learning with Reduced Information Leakage and Computation

### v1

@@ Abstract
Federated learning (FL) is a distributed learning paradigm that allows multiple decentralized clients to collaboratively learn a common model without sharing local data. Although local data is not exposed directly, privacy concerns nonetheless exist as clients’ sensitive information can be inferred from intermediate computations. Moreover, such information leakage accumulates substantially over time as the same data is repeatedly used during the iterative learning process. As a result, it can be particularly difficult to balance the privacy-accuracy trade-off when designing privacy-preserving FL algorithms. In this paper, we introduce Upcycled-FL, a novel federated learning framework with first-order approximation applied at every even iteration. Under this framework, half of the FL updates incur no information leakage and require much less computation. We first conduct the theoretical analysis on the convergence (rate) of Upcycled-FL, and then apply perturbation mechanisms to preserve privacy. Experiments on real-world data show that Upcycled-FL consistently outperforms existing methods over heterogeneous data, and significantly improves privacy-accuracy trade-off while reducing 48% of the training time on average.
@@ 1 Introduction
Federated learning has emerged as an important paradigm for learning models in a distributed fashion, whereby data is distributed across different clients and the goal is to jointly learn a model from the distributed data. This is facilitated by a central server and the model can be learned through iterative interactions between the central server and clients: at each iteration, each client performs certain computation using its local data; the local results are collected and aggregated by the central server; the aggregated result is then sent to clients to update local results; and so on till the learning task is deemed accomplished.
Although each client’s data is not shared directly with the central server, their sensitive information may be inferred from the intermediate computations. Indeed, as the same data is repeatedly used during the iterative learning process, each client’s total information leakage accumulates substantially over time. For randomization-based privacy-preserving techniques such as differential privacy [1], the perturbation required for guaranteeing the privacy of each client also grows as the total information leakage increases. As a result, it can be particularly difficult to balance the privacy and accuracy trade-off in FL.
@@
This paper proposes a novel federated learning framework called Upcycled Federated Learning (Upcycled-FL)***The word “upcycle” refers to reusing material so as to create higher-quality things than the original., in which clients’ information leakage can be reduced such that it only occurs during half of the updates. This is attained by modifying the even iterations of a baseline FL algorithm with first-order approximation, which allows us to compute the update using existing results from previous iterations without using the client’s data. Moreover, the updates in even iterations only involve addition/subtraction operations on existing results at the central server; the computational costs and the training time can be reduced significantly.
To preserve privacy, we can apply existing privacy mechanisms to Upcycled-FL. As an example, we consider differential privacy (DP) because it can (i) rigorously quantify the total privacy leakage for complex algorithms such as FL; (ii) defend against attackers regardless of their background knowledge; (iii) provide heterogeneous privacy guarantees for different clients. Since DP preserves the client’s privacy by randomizing the algorithm (e.g., adding noise to the output, objective, or gradient of local clients), Upcycled-FL by reducing the total information leakage, requires less randomization to attain a certain level of privacy. Therefore, the privacy-accuracy trade-off is also improved.
@@
We emphasize that the idea of upcycling is orthogonal to both the baseline FL algorithm, and the DP perturbation method. In this paper, we adopt FedProx [2] as the baseline FL algorithm and apply first-order approximation to it, but our approach can be applied to other FL frameworks, such as FedNova [3] and pFedMe [4]. For DP perturbation methods, we consider both output and objective perturbation as examples to quantify the privacy loss, while other DP methods can also be used.
@@
It is worth noting that although differentially private federated learning has been extensively studied in the literature, e.g., [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from this work, where we propose a novel method for FL that reduces information leakage and computation.
@@
In addition to private federated learning, many approaches were proposed in the DP literature to improve the privacy-accuracy trade-off (e.g., privacy amplification by sampling [21, 22, 23, 24, 25, 26, 27], leveraging non-private public data [28, 29], shuffling [30], using weaker privacy notion [31], using tighter privacy composition analysis tool [27]). However, none of these strategies affect the algorithmic properties of learning algorithms. By contrast, our method improves the privacy-accuracy trade-off by modifying the property of FL algorithm (i.e., reducing the total information leakage); this improvement on the algorithmic property is independent of the privacy notion/mechanism or the analysis method. More related work is discussed in Appendix B
@@
  • We propose Upcycled-FL (Algorithm 2), a novel federated learning framework with reduced information leakage and computation.
  • We conduct convergence (rate) analysis (Section 4, Theorem 4.6) and identify a sufficient condition for the convergence of Upcycled-FL.
  • We apply output perturbation and objective perturbation to Upcycled-FL and conduct privacy analysis (Section 5, Theorem C.1).
  • We evaluate the performance of (private) Upcycled-FL on real-world data (Section 6). The results show that Upcycled-FL, by reducing information leakage, outperforms existing algorithms over heterogeneous data and improves the accuracy-privacy trade-off.
2 Problem Formulation and Preliminaries
Consider an FL system consisting of a central server and a set $\mathcal{I}$ of clients. Each client $i$ has its local data $\mathcal{D}_{i}$ and they can be non-i.i.d across the clients. The goal of FL is to learn a model $\omega\in\mathbb{R}^{d}$ from data $\cup_{i\in\mathcal{I}}\mathcal{D}_{i}$ by solving the following optimization:
@@ 2 Problem Formulation and Preliminaries
FL Algorithm. Let $\omega_{i}^{t}$ be client $i$’s local model parameter at time $t$. In FL, the model is learned through an iterative process: at each time step $t$, (1) local computations: each active client updates its local model $\omega_{i}^{t}$ using its local data $\mathcal{D}_{i}$; (2) local models broadcasts: local models (or gradients) are then uploaded to the central server; (3) model aggregation: the central server aggregates results received from clients to update the global model parameter $\overline{\omega}^{t}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$; (4) model updating: the aggregated model is sent back to clients and is used for updating local models at $t+1$. In this work, we consider a baseline FL framework FedProx and apply the proposed method to it, while our method can also be used on other FL frameworks such as FedNova [3] and pFedMe [4]. The details of FedProx are given in Algorithm 1.
Algorithm 1 FedProx [2]
1:  Input: $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
2:  for $t=1$ to $T$ do
5:     Each client sends its local model to server.
$\overline{\omega}^{t+1}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t+1}.$
7:  end for
@@
Differential Privacy [1]. Differential privacy (DP) centers around the idea that the output of a certain computational procedure should be statistically similar given singular changes to the input, thereby preventing meaningful inference from observing the output.
In FL, the information exposed by each client $i$ includes all intermediate computations $\{\omega_{i}^{t}\}_{t=1}^{T}$. Consider a randomized FL algorithm $\mathcal{A}(\cdot)$ that generates a sequence of private local models $\{\widehat{\omega}_{i}^{t}\}_{t=1}^{T}$, we say it satisfies DP for client $i$ over $T$ iterations if the following holds for any possible output $O\in\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}$, and for any two neighboring local datasets $\mathcal{D}_{i}$, $\mathcal{D}^{\prime}_{i}$:
@@
3 Proposed Algorithm: Upcycled-FL
@@
Main idea: Fundamentally, the accumulation of information leakage over iterations stems from the fact that the client’s data $\mathcal{D}_{i}$ is used in every update. If the updates can be made without directly using this original data, but only from computational results that already exist, then the information leakage originating from these updates will be zero, and meanwhile the computational cost may be reduced significantly. Based on this idea, we propose Upcycled-FL, which modifies FedProx such that the earlier computations are repeatedly used for new updates.
@@
Upcycling Information: Next, we present Upcycled-FL and illustrate how the client’s total information leakage is reduced under this method.
@@
We modify FedProx by applying first-order approximation to even iterations. Specifically, at even iteration $2m$, we expand $F_{i}(\omega;\mathcal{D}_{i})$ at $\omega^{2m-1}_{i}$ (local model in the previous iteration). Based on Taylor series expansion, we have:
$F_{i}(\omega;\mathcal{D}_{i})$ $=$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$
$+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})$
$+\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2}).$
We approximate term $\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2})$ by $\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}$ for some constant $\lambda_{m}\geq 0$, which can be different for different iteration $2m$, then the even update of FedProx under such an approximation becomes:
@@
$\omega_{i}^{2m}$ $=$ $\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$ (2)
$\approx$ $\arg\min_{\omega}F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}$
$=$ $\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega$
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
@@
Note that in the above even update, the only term that depends on dataset $\mathcal{D}_{i}$ is $\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$, which can be derived directly from the previous odd iteration. Specifically, according to the first-order condition, the following holds at odd iterations:
$\omega_{i}^{2m-1}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2}$
@@
Plug $\nabla F_{i}(\omega_{i}^{2m-1};\mathcal{D}_{i})$ into even updates (2), we have
$\omega_{i}^{2m}$$\approx$ $\arg\min_{\omega}\mu(\overline{\omega}^{2m-2}-\omega_{i}^{2m-1})^{T}\omega$ (4)
$+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$
By first-order condition, even updates (4) can be reduced to:
@@
It turns out that with first-order approximation, dataset $\mathcal{D}_{i}$ is not used in the even updates. Furthermore, these updates can be conducted at the central server directly. That is, the central server updates the global model by aggregating:
$\overline{\omega}^{2m}$ $=$ $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m}$
$\approx$ $\sum_{i\in\mathcal{I}}p_{i}\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
$=$ $\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
Therefore, even updates only involve addition/subtraction operations on the existing global models from previous iterations (i.e., $\overline{\omega}^{2m-1},\overline{\omega}^{2m-2}$): the computational cost and training time is reduced significantly. Note that the first-order approximation is only applied to even iterations, while the odd iterations should remain the same as FedProx to ensure Eqn. (3) holds. The entire updating procedure of Upcycled-FL is summarized in Algorithm 2.
@@
Discussion: Indeed, if we view two consecutive steps (odd $12m-$ 1 followed by even $2m$) of Upcycled-FL as a single iteration $t$, then Upcycled-FL and FedProx will incur the same information leakage but will differ at the phase of global model aggregation. Specifically, instead of simply aggregating the global model by averaging over local updates (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$) as in FedProx, Upcycled-FL not only takes the average of local updates but also push the aggregation moving more toward the updating direction (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}-\overline{\omega}^{t-1}$). In Figure 1, we illustrate the difference between FedProx and Upcycled-FL from the two perspectives: (i) Upcycled-FL can be regarded as reusing the intermediate updates of FedProx, compared to which the total information leakage is reduced; or (ii) Upcycled-FL can be regarded as FedProx with a different global aggregation rule, and the larger global update in Upcycled-FL accelerates the learning process with the same information leakage.
@@
Figure 1: Consider Upcycled-FL from two perspectives

### v2

# algorithmic

@@ Abstract
Federated learning (FL) is a distributed learning paradigm that allows multiple decentralized clients to collaboratively learn a common model without sharing local data. Although local data is not exposed directly, privacy concerns nonetheless exist as clients’ sensitive information can be inferred from intermediate computations. Moreover, such information leakage accumulates substantially over time as the same data is repeatedly used during the iterative learning process. As a result, it can be particularly difficult to balance the privacy-accuracy trade-off when designing privacy-preserving FL algorithms. This paper introduces Upcycled-FL, a simple yet effective strategy that applies first-order approximation at every even round of model update. Under this strategy, half of the FL updates incur no information leakage and require much less computational and transmission costs. We first conduct the theoretical analysis on the convergence (rate) of Upcycled-FL and then apply two perturbation mechanisms to preserve privacy. Extensive experiments on both synthetic and real-world data show that the Upcycled-FL strategy can be adapted to many existing FL frameworks and consistently improve the privacy-accuracy trade-off.*** Code available at https://github.com/osu-srml/Upcycled-FL.
@@ 1 Introduction
Federated learning (FL) has emerged as an important paradigm for learning models in a distributed fashion, whereby data is distributed across different clients and the goal is to jointly learn a model from the distributed data. This is facilitated by a central server and the model can be learned through iterative interactions between the central server and clients: at each iteration, each client performs certain computation using its local data; the updated local models are collected and aggregated by the server; the aggregated model is then sent back to clients for them to update local models; and so on till the learning task is deemed accomplished.
Although each client’s data is not shared directly with the central server, there is a risk of information leakage that a third party may infer individual sensitive information from (intermediate) computational outcomes. Information leakage occurs whenever the client’s local gradients are shared with third parties. Importantly, the information leakage (or privacy loss) accumulates as data is repeatedly used during the iterative learning process: with more computational outcomes derived from individual data, third parties have more information to infer sensitive data and it poses higher privacy risks for individuals. An example is Huang et al. (2021), which shows that eavesdroppers can conduct gradient inversion attacks to recover clients’ data from the gradients. In this paper, we use differential privacy (DP) proposed by Dwork (2006), a de facto standard for preserving individual data privacy in data analysis, ranging from simple tasks such as data collection and statistical analysis (Zhang et al., 2022b; Ghosh and Roth, 2011; Khalili and Vakilinia, 2021; Amin et al., 2019; Liu et al., 2021; Khalili et al., 2021c, 2019) to complex machine learning and optimization tasks (Cai et al., 2024; Khalili et al., 2021b, a; Huang et al., 2020; Jagielski et al., 2019; Zhang et al., 2019b, 2018a, 2018b, c). Compared to other privacy preservation techniques, it can (i) rigorously quantify the total privacy leakage for complex algorithms such as FL; (ii) defend against attackers regardless of their background knowledge; (iii) provide heterogeneous privacy guarantees for different clients. To achieve a certain DP guarantee, we need to perturb FL algorithms (e.g., adding noise to the output, objective, or gradient of local clients) and the perturbation needed for achieving the privacy requirement of each client grows as the total information leakage increases. Because the added perturbation inevitably reduces algorithm accuracy, it can be difficult to balance the privacy and accuracy trade-off in FL.
@@
This paper proposes a novel strategy for FL called Upcycled Federated Learning (Upcycled-FL)The word “upcycle” refers to reusing material so as to create higher-quality things than the original., in which clients’ information leakage can be reduced such that it only occurs during half of the updates. This is attained by modifying the even iterations of a baseline FL algorithm with first-order approximation, which allows us to update the model using existing model parameters from previous iterations without using the client’s data. Moreover, the updates in even iterations only involve addition/subtraction operations on existing model parameters at the central server. Because Upcycled-FL doesn’t require local training and transmission in half of iterations, the transmission costs and the training time can be reduced significantly. It turns out that Upcycled-FL, by reducing the total information leakage, requires less perturbation to attain a certain level of privacy and can enhance the privacy-accuracy trade-off significantly.
@@
We emphasize that the idea of “upcycling information” is orthogonal to both the baseline FL algorithm and the DP perturbation method. It can be applied to any FL algorithms that involve local optimization at the clients. In this paper, we apply Upcycled-FL strategy to multiple existing FL algorithms and evaluate them on both synthetic and real-world datasets. For DP perturbation methods, we consider both output and objective perturbation as examples to quantify the privacy loss, while other DP methods can also be used.
@@
It is worth noting that although differentially private federated learning has been extensively studied in the literature, e.g., (Asoodeh et al., 2021; Chuanxin et al., 2020; Zhang et al., 2022a; Zheng et al., 2021; Wang et al., 2020b; Kim et al., 2021; Zhang et al., 2019a; Baek et al., 2021; Wu et al., 2022; Girgis et al., 2021; Truex et al., 2020; Hu et al., 2020; Seif et al., 2020; Zhao et al., 2020; Wei et al., 2020; Triastcyn and Faltings, 2019), all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from this work, where we propose a novel strategy that effectively reduces information leakage in FL.
@@
In addition to private federated learning, several approaches were proposed in the DP literature to improve the privacy-accuracy trade-off, e.g., privacy amplification by sampling (Balle et al., 2018; Beimel et al., 2014; Hu et al., 2021; Wang et al., 2019; Kasiviswanathan et al., 2011; Wang et al., 2015; Abadi et al., 2016), leveraging non-private public data (Avent et al., 2017; Papernot et al., 2016), shuffling (Úlfar Erlingsson et al., 2019), using weaker privacy notion (Bun and Steinke, 2016), using tighter privacy composition analysis tool (Abadi et al., 2016). However, none of these strategies affect the algorithmic properties of learning algorithms. By contrast, our method improves the privacy-accuracy trade-off by modifying the property of FL algorithm (i.e., reducing the total information leakage); this improvement on the algorithmic property is independent of the privacy notion/mechanism or the analysis method.
@@
  • We propose Upcycled-FL (Algorithm 1), a novel strategy with reduced information leakage and computation that is broadly applicable to many existing FL algorithms.
  • As an example, we apply our strategy to FedProx (Li et al., 2020) and conduct convergence (rate) analysis (Section 5, Theorem 5.6) where we identify a sufficient condition for the convergence of Upcycled-FL.
  • As an example, we apply two differential privacy mechanisms (i.e., output perturbation and objective perturbation) to Upcycled-FL and conduct privacy analysis (Section 6, Theorem 6.2).
  • We evaluate the effectiveness of Upcycled-FL on both synthetic and real data (Section 7). Extensive experiments show that Upcycled-FL can be adapted to many existing federated algorithms to achieve better performance; it effectively improves the accuracy-privacy trade-off by reducing information leakage.
2 Related Work
Differential privacy in federated learning. Differential privacy has been widely used in federated learning to provide privacy guarantees (Asoodeh et al., 2021; Chuanxin et al., 2020; Zhang et al., 2022a; Zheng et al., 2021; Wang et al., 2020b; Kim et al., 2021; Zhang et al., 2019a; Baek et al., 2021; Wu et al., 2022; Girgis et al., 2021; Truex et al., 2020; Hu et al., 2020; Seif et al., 2020; Zhao et al., 2020; Wei et al., 2020; Triastcyn and Faltings, 2019). For example, Zhang et al. (2022a) uses the Gaussian mechanism for a federated learning problem and propose an incentive mechanism to encourage users to share their data and participate in the training process. Zheng et al. (2021) introduces $f$-differential privacy, a generalized version of Gaussian differential privacy, and propose a federated learning algorithm satisfying this new notion. Wang et al. (2020b) proposes a new mechanism called Random Response with Priori (RRP) to achieve local differential privacy and apply this mechanism to the text data by training a Latent Dirichlet Allocation (LDA) model using a federated learning algorithm. Triastcyn and Faltings (2019) adapts the Bayesian privacy accounting method to the federated setting and propose joint accounting method for estimating client-level and instance-level privacy simultaneously and securely. Wei et al. (2020) presents a private scheme that adds noise to parameters at the random selected devices before aggregating and provides a convergence bound. Kim et al. (2021) combines the Gaussian mechanism with gradient clipping in federated learning to improve the privacy-accuracy tradeoff. Asoodeh et al. (2021) considers a different setting where only the last update is publicly released and the central server and other devices are assumed to be trustworthy. However, all these algorithms need client’s local data to update the model and the information leakage inevitably occurs at every iteration. This is fundamentally different from Upcycled-FL, which reuses existing results to update half of iterations and significantly reduces information leakage and computation.
Tackling heterogeneity in federated learning. It’s worth mentioning that, Upcycled-FL also empirically outperforms existing baseline algorithms under device and statistical heterogeneity. In real-world scenarios, local data are often non-identically distributed across different devices; different devices are also often equipped with different specifications and computation capabilities. Such heterogeneity often causes instability in the model performance and leads to divergence. Many approaches have been proposed to tackle this issue in FL. For example, FedAvg (McMahan et al., 2017) uses a random selection of devices at each iteration to reduce the negative impact of statistical heterogeneity; however, it may fail to converge when heterogeneity increases. Other methods include FedProx (Li et al., 2020), a generalization and re-parameterization of FedAvg that adds a proximal term to the objective function to penalize deviations in the local model from the previous aggregation, and FedNova (Wang et al., 2020a) that re-normalizes local updates before updating to eliminate objective inconsistency. It turns out that Upcycled-FL exhibits superior performance in the presence of heterogeneity because gradients encapsulate information on data heterogeneity, the reusing of which leads to a boost in performance.
3 Problem Formulation
Consider an FL system consisting of a central server and a set $\mathcal{I}$ of clients. Each client $i$ has its local dataset $\mathcal{D}_{i}$ and these datasets can be non-i.i.d across the clients. The goal of FL is to learn a model $\omega\in\mathbb{R}^{d}$ from data $\cup_{i\in\mathcal{I}}\mathcal{D}_{i}$ by solving the following optimization:
@@ 3 Problem Formulation
FL Algorithm. Let $\omega_{i}^{t}$ be client $i$’s local model parameter at time $t$. In FL, the model is learned through an iterative process: at each time step $t$, 1) local computations: each active client updates its local model $\omega_{i}^{t}$ using its local data $\mathcal{D}_{i}$; 2) local models broadcasts: local models (or gradients) are then uploaded to the central server; 3) model aggregation: the central server aggregates results received from clients to update the global model parameter $\overline{\omega}^{t}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$; 4) model updating: the aggregated model is sent back to clients and is used for updating local models at $t+1$.
@@
Differential Privacy (Dwork, 2006). Differential privacy (DP) centers around the idea that the output of a certain computational procedure should be statistically similar given singular changes to the input, thereby preventing meaningful inference from observing the output.
In FL, the information exposed by each client $i$ includes all intermediate computations $\{\omega_{i}^{t}\}_{t=1}^{T}$. Consider a randomized FL algorithm $\mathcal{A}(\cdot)$ that generates a sequence of private local models $\{\widehat{\omega}_{i}^{t}\}_{t=1}^{T}$, we say it satisfies $(\varepsilon,\delta)$-differential privacy for client $i$ over $T$ iterations if the following holds for any possible output $O\in\mathbb{R}^{d}\times\cdots\times\mathbb{R}^{d}$, and for any two neighboring local datasets $\mathcal{D}_{i}$, $\mathcal{D}^{\prime}_{i}$:
@@
4 Proposed Method: Upcycled-FL
Main idea.
@@ 4 Proposed Method: Upcycled-FL
Fundamentally, the accumulation of information leakage over iterations stems from the fact that the client’s data $\mathcal{D}_{i}$ is used in every update. If the updates can be made without directly using this original data, but only from computational results that already exist, then the information leakage originating from these updates will be zero, and meanwhile, the computational cost may be reduced significantly. Based on this idea, we propose Upcycled-FL, which considers reusing the earlier computations in a new update and significantly reduces total information leakage and computational cost. Note that Upcycled-FL is not a specific algorithm but an effective strategy that can be used for any existing FL algorithms.
Upcycling model update.
@@
Next, we present Upcycled-FL and illustrate how the client’s total information leakage is reduced under this method.
@@
For an FL system with the objective shown in Eqn. (1), the client $i$’s local objective function is given by $F_{i}(\omega;\mathcal{D}_{i})$. Under Upcycled-FL, we apply first-order approximation to $F_{i}(\omega;\mathcal{D}_{i})$ at even iterations during federated training (while odd updates remain fixed). Specifically, at $2m$-th iteration, we expand $F_{i}(\omega;\mathcal{D}_{i})$ at $\omega^{2m-1}_{i}$ (local model in the previous iteration). Based on the Taylor series expansion, we have:
$F_{i}(\omega;\mathcal{D}_{i})$ $=$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})+\mathcal{O}(||\omega-\omega^{2m-1}_{i}||^{2})$ (2)
$\approx$ $F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})+\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}(\omega-\omega^{2m-1}_{i})+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}$
for some constant $\lambda_{m}\geq 0$ which may differ for different iteration $2m$. Then for an FL algorithm, its model update at $2m$-th iteration under Upcycled-FL strategy can be attained by replacing $F_{i}(\omega;\mathcal{D}_{i})$ with its approximation in Eqn. (2) while the updates at odd iterations remain the same. We illustrate this using the following two examples.
Example 4.1 (FedAvg (McMahan et al., 2017) under Upcycled-FL strategy).
In FedAvg, client $i$ at each iteration updates the local model by optimizing its local objective function, i.e., $\omega_{i}^{t}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i}),\forall t$. Under Upcycled-FL strategy, the client $i$’s updates become:
$\omega_{i}^{t}=\begin{cases}\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2},~{}~{}\text{ if }~{}~{}t=2m\\ \arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i}),~{}~{}\text{ if }~{}~{}t=2m-1\par\end{cases}$
Example 4.2 (FedProx (Li et al., 2020) under Upcycled-FL strategy).
In FedProx, a proximal term is added to the local objective function to stabilize the algorithm under heterogeneous clients (Algorithm 2), i.e., client $i$ at each iteration updates the local model $\omega_{i}^{t}=\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{t-1}||^{2},\forall t$. Under Upcycled-FL strategy, the client $i$’s updates become:
@@
$\omega_{i}^{t}=\begin{cases}\arg\min_{\omega}\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2},~{}~{}\text{ if }~{}~{}t=2m\\ \arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-2}||^{2},~{}~{}\text{ if }~{}~{}t=2m-1\end{cases}$ (3)
Next, we demonstrate how the information is upcycled under the above idea. As an illustrating example, we focus on FedProx given in Example 4.2.
@@
Note that in the even update of Eqn. (3), the only term that depends on dataset $\mathcal{D}_{i}$ is $\nabla F_{i}(\omega^{2m-1}_{i};\mathcal{D}_{i})$, which can be derived directly from the previous odd iteration. Specifically, according to the first-order condition, the following holds at odd iterations:
@@
Plug $\nabla F_{i}(\omega_{i}^{2m-1};\mathcal{D}_{i})$ into even update of (3), we have the estimated update from the odd update:
$\omega_{i}^{2m}$$\approx$ $\arg\min_{\omega}\mu(\overline{\omega}^{2m-2}-\omega_{i}^{2m-1})^{T}\omega+\frac{\lambda_{m}}{2}||\omega-\omega^{2m-1}_{i}||^{2}+\frac{\mu}{2}||\omega-\overline{\omega}^{2m-1}||^{2}.$ (5)
where the approximately equals sign $\approx$” is due to the approximation in Eqn. (2). By first-order condition, even update (5) can be reduced to:
@@
It turns out that with first-order approximation, dataset $\mathcal{D}_{i}$ is not used in the even updates. Because these updates do not require access to the client’s data, these updates can be conducted at the central server directly. That is, the central server updates the global model by aggregating:
$\overline{\omega}^{2m}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{2m}\approx\sum_{i\in\mathcal{I}}p_{i}\omega^{2m-1}_{i}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)=\overline{\omega}^{2m-1}+\frac{\mu}{\mu+\lambda_{m}}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
Therefore, under Upcycled-FL strategy, even updates only involve addition/subtraction operations on the existing global models from previous iterations (i.e., $\overline{\omega}^{2m-1},\overline{\omega}^{2m-2}$) without the need for a local training epoch: both the computational cost and transmission cost are reduced significantly. Note that the first-order approximation is only applied to even iterations, while the odd iterations should remain as the origin to ensure Eqn. (4) holds. The entire updating procedure of Upcycled-FL is summarized in Algorithm 1.
@@
Figure 1: Upcycled-FL can be considered from two perspectives: (i) it can be regarded as reusing the intermediate updates of local models to reduce the total information leakage; or (ii) it can be regarded as a global aggregation method with larger global update, which accelerates the learning process with the same information leakage under the same training iterations.
Discussion.
Indeed, if we view two consecutive steps (odd $12m-$ 1 followed by even $2m$) of Upcycled-FL as a single iteration $t$, then Upcycled-FedProx and FedProx will incur the same information leakage but will differ at the phase of global model aggregation, as shown in Figure 1(b). Specifically, instead of simply aggregating the global model by averaging over local updates (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}$), Upcycled-FL not only takes the average of local updates but also pushes the aggregation moving more toward the updating direction (i.e., $\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t}-\overline{\omega}^{t-1}$). We present the difference between Upcycled-FL update strategy and regular aggregation strategy from these two perspectives in Figure 1. As Upcycled-FL only accesses client data at even iterations, it halves the communication cost compared to standard FL methods with the same number of iterations.
Algorithm 1 Proposed aggregation strategy: Upcycled-FL
2:  for $m=1$ to $M$ do
3:    The central server sends the global model parameters $\overline{\omega}^{2m-2}$ to all the clients.
4:    A subset of clients are selected to be active and each active client updates its local model by finding an (approximate) minimizer of the local objective function:
$\omega_{i}^{2m-1}\leftarrow\arg\min_{\omega}F_{i}(\omega;\mathcal{D}_{i})~{}~{}~{}~{}\text{ or~{} { other local objective functions}}$
5:    Clients send local models to the central server.
$\overline{\omega}^{2m}$ $=$ $\overline{\omega}^{2m-1}+\frac{\lambda_{m}}{2}\left(\overline{\omega}^{2m-1}-\overline{\omega}^{2m-2}\right)$
7:  end for